In cucumbers, dull fruit (D) is dominant over glossy fruit (d), orange fruit (R) is dominant over cream fruit (r), and bitter cotyledons (B) are dominant over non-bitter cotyledons (b). The three characters are encoded by genes located on different pairs of chromosomes. A plant homozygous for dull, orange fruit and bitter cotyledons is crossed with a plant that has glossy, cream fruit and non-bitter cotyledons. The F1 are intercrossed to produce the F2.

a. Give the phenotypes and their expected proportions in the F2.

b. An F1 plant is crossed with a plant that has glossy, cream fruit and non-bitter cotyledons. Give the phenotypes and expected proportions among the progeny of this cross.
Answer #1

The genotype of dull, orange fruit and bitter cotyledons = DDRRBB

The genotype of glossy, cream fruit and non-bitter = ddrrbb

In f1 generation, all the progeny will have DdRrBb genotype.

In the f2 generation number of phenotypes can be calculated by the probability method.

First, we calculate the probability for a monohybrid cross and then phenotypes fraction obtained by combining them

Dull fruit Dd x Dd

D d
D DD Dd
d Dd dd

Probability of the dull phenotype = 3/4

Probability of the glossy phenotype = 1/4

Orange fruit Rr x Rr

R r
R RR Rr
r Rr rr

Probability of the Orange fruit = 3/4

Probability of the Cream fruit = 1/4

Bitter fruit Bb x Bb

B b
B BB Bb
b Bb bb

​​​probability of Bitter cotyledons = 3/4, probability of nonbitter cotyledon = 1/4